# Digital-physical testbed for ship autonomy studies in the Marine Cybernetics Laboratory basin

### v1

@@ 5 Conclusions
While the digital-physical testbed provides a valuable platform for development and testing, it has some limitations. Small-scale models does not fully reflect full-scale behavior, and simulations can only approximate real-world conditions. The controlled lab environment also limits exposure to the variability of actual maritime operations. Finally, the resource demands of the testbed may affect its scalability.
@@ 6 Acknowledgments
Work funded by the Research Council of Norway (RCN) through SFI AutoShip(RCN project 309230), NTNU AMOS (RCN project 223254), SFI Harvest (RCN project 309661), and Partly funded by the European Union via the LORELEI-X Twinning project (GA 101159489). Authors would also like to thank Jon Estil Krågebakk for providing 3D models, and the team of engineers at MC-Lab, Robert Opland and Vebjørn Steinsholt.

### v2

@@ 5 Conclusions
While the digital-physical testbed offers a valuable platform for development and testing, it has certain limitations. First, small-scale models do not fully capture the behavior of full-scale systems, and simulations can only approximate real-world conditions. Moreover, a bollard-pull test—which is commonly used in marine engineering to measure the thrust generated by a vessel’s propulsion system at rest—has not yet been conducted, limiting the accuracy of the thrust modeling. The controlled laboratory environment also restricts exposure to the variability and unpredictability of real maritime operations. Lastly, the testbed’s resource requirements may pose challenges to its scalability.
@@ 6 Acknowledgments
Work funded by the Research Council of Norway (RCN) through SFI AutoShip(RCN project 309230), NTNU AMOS (RCN project 223254), SFI Harvest (RCN project 309661), and partly funded by the European Union via the LORELEI-X Twinning project (GA 101159489) and NTNU VISTA CAROS. Authors would also like to thank Jon Estil Krågebakk for providing 3D models, and the team of engineers at MC-Lab, Robert Opland, and Vebjørn Steinsholt.